OpenStudy (anonymous):

if a=8, b=6, and c=6 then what are the angles of A,B,and C in degrees?

OpenStudy (anonymous):

use cosine rule,cosA=(b^2+c^2-a^2)/2bc

OpenStudy (radar):

 \[\cos A=6^{2} + 6^{2}- 8^{2}/((2)(6)(6)) =8/72=.111111\] Angle A=83.62 degrees. The other two angles are equal to each other (isosceles triangle) 2x+83.62 = 180 degrees 2x=96.38 x=48.19, Angle b, angle c are 48.19 degrees each.